# LoXR: Performance Evaluation of Locally Executing LLMs on XR Devices

### v1

@@ 6 Discussion
Limitations: Apart from the AVP, our study primarily focuses on CPU-based evaluation and does not fully explore the potential of GPU- and NPU-accelerated inference for on-device LLM deployment. The limited VRAM on the three devices (ML2, Vivo, and MQ3) restricted full model loading, which hindered GPU performance. Additionally, the specialized frameworks and configurations required for GPU and NPU optimization were beyond the scope of this work. Finally, since the study is primarily based on Llama.cpp implementations, its generalizability is limited.
Future Research: In future research, we plan to conduct an in-depth exploration of GPU and NPU-based deployment, aiming to optimize performance for lower-memory GPUs while maximizing efficiency. Furthermore, the integration of visual language models (VLMs) capable of processing images and graphical inputs will be ideally suited for the XR domain, expanding potential applications across various fields, including education, healthcare, and robotics. Additionally, voice- or text-based interaction with on-device LLMs can lead to several XR applications. For example, training the voice system with a specific teacher’s voice and utilizing it as a communication medium for a teacher’s virtual avatar could create a compelling educational tool—a virtual tutor. A similar approach could also be applied to humanoid robots (or avatars in VR), enabling more natural and interactive communication.
@@ 7 Conclusion
LoXR presents a comprehensive study on deploying LLMs on four XR devices, conducting an in-depth analysis of multiple performance factors. Given the GPU memory limitations of most XR devices, our study primarily focuses on CPU-based analysis. However, since the AVP has sufficient GPU resources, we also include GPU results for AVP, providing additional insights into the study. Generally, for processing speed, we can rank the four devices in the following order: AVP, ML2, Vivo, and MQ3, although MQ3 outperforms Vivo in TG.
In memory consumption, AVP was found the most memory-efficient. Among the remaining devices, differences were minimal, though ML2 occasionally consumed the most memory. As expected, larger models required more memory, but this trend varied across different model series. For BT tests, a lower batch size (128) yielded the best performance, while larger batch sizes reduced speed for CPU-based tests, with AVP GPU remaining largely unaffected. For TT tests, medium thread counts (4–8) were optimal across all devices. In terms of consistency, ML2 is the most stable, exhibiting the lowest variance in speed and zero errors. AVP GPU also maintains low variance. MQ3 performs the worst in this aspect, while Vivo is more stable than MQ3 but less consistent than the rest.
The Pareto analysis highlighted several models on AVP due to their outstanding performance, while the Pareto fronts found on ML2 because of their consistently stable results. Additionally, a few models on Vivo and MQ3 emerged on the Pareto front because of their balanced trade-offs between performance and stability. Models $m_{12}$ to $m_{17}$ were significantly slower.

### v2

@@ 7.6 Limitations
Apart from AVP, our study focuses on CPU-based evaluation, as limited GPU memory on the three devices hindered full model loading, affecting performance. Finally, since the study is primarily based on Llama.cpp implementations, its generalizability is limited.
@@ 8 Conclusion
AIvaluateXR provides a comprehensive framework for deploying LLMs on XR devices and conducting in-depth evaluations across multiple performance factors. Using this framework, we benchmarked 17 LLMs across four XR devices to assess their performance in a systematic and reproducible manner. Given the GPU memory limitations of most XR devices, our study primarily focuses on CPU-based analysis. However, since the AVP has sufficient GPU resources, we also include GPU results for AVP, providing additional insights into the study. Generally, for processing speed, we can rank the four devices in the following order: AVP, ML2, Vivo, and MQ3, although MQ3 outperforms Vivo in TG.
In memory consumption, AVP was found the most memory-efficient. Among the remaining devices, differences were minimal, though ML2 occasionally consumed the most memory. As expected, larger models required more memory, but this trend varied across different model series. For BT tests, a lower batch size (128) yielded the best performance, while larger batch sizes reduced speed for CPU-based tests, with AVP GPU remaining largely unaffected. For TT tests, medium thread counts (4–8) were optimal across all devices. In terms of consistency, ML2 is the most stable, exhibiting the lowest variance in speed and zero errors. AVP GPU also maintains low variance. MQ3 performs the worst in this aspect, while Vivo is more stable than MQ3 but less consistent than the rest. The Pareto analysis highlighted several models on AVP due to their outstanding performance, while the Pareto fronts found on ML2 because of their consistently stable results. Additionally, a few models on Vivo and MQ3 emerged on the Pareto front because of their balanced trade-offs between performance and stability. Models $m_{12}$ to $m_{17}$ were found slower.
Our latency comparison across cloud-based, client-server, and on-device deployments—along with two interactive application tests—demonstrates that while on-device LLMs currently yield slightly lower accuracy, this can be improved through domain-specific fine-tuning. Notably, despite longer warm-up times (i.e., higher Cold F.R.T), on-device models offer excellent latency post-loading (i.e., lower Warm F.R.T), making them highly promising for real-time XR applications. Given the rise of XR and LLMs, we believe on-device AI has strong potential, in which AIvaluateXR can play a key role.
Acknowledgments
This research has been funded by KAUST Competitive Research Grants ORFS-CRG12-2024-6422. We thank Prof. Kiyoshi Kiyokawa from NAIST, Japan, for valuable discussions. We also thank Deng Luo and Da Li from our team at KAUST for their support and valuable input across various modules.